# Multi-User Chat Assistant (MUCA): a Framework Using LLMs to Facilitate Group Conversations

### 10 Jan 2024 v1

@@ Abstract
Recent advancements in large language models (LLMs) have provided a new avenue for chatbot development, while most existing research has primarily centered on single-user chatbots that focus on deciding "What" to answer after user inputs. In this paper, we identified that multi-user chatbots have more complex 3W design dimensions – "What" to say, "When" to respond, and "Who" to answer. Additionally, we proposed Multi-User Chat Assistant (MUCA), which is an LLM-based framework for chatbots specifically designed for group discussions. MUCA consists of three main modules: Sub-topic Generator, Dialog Analyzer, and Utterance Strategies Arbitrator. These modules jointly determine suitable response contents, timings, and the appropriate recipients. To make the optimizing process for MUCA easier, we further propose an LLM-based Multi-User Simulator (MUS) that can mimic real user behavior. This enables faster simulation of a conversation between the chatbot and simulated users, making the early development of the chatbot framework much more efficient. MUCA demonstrates effectiveness, including appropriate chime-in timing, relevant content, and positive user engagement, in goal-oriented conversations with a small to medium number of participants, as evidenced by case studies and experimental results from user studies.
@@ 2 Related Work
Open-Domain Models: LLMs have demonstrated remarkable performance across various tasks. They are typically pre-trained on large amounts of data and fine-tuned with human-curated data through supervised learning and reinforcement learning approaches. Some of the recent work including GPT models OpenAI (2023); Ouyang et al. (2022); Brown et al. (2020); Radford et al. (2019b), LLaMa2 Touvron et al. (2023), PaLM2 Anil et al. (2023), Bloom Scao et al. (2022) and LaMDa Thoppilan et al. (2022). Some of the recent versions of these models, such as GPT-4, have shown promising results as conversational agents. Despite the overall good performance of these models across multiple tasks on text generation, there are still continuing efforts to improve them for specific scenarios Chen et al. (2021); Wu et al. (2023b).
@@ 3.2 Framework Architecture
Figure 3 illustrates the framework architecture and information flow, which consists of three major modules: (1) Sub-topics Generator, which generates the initial sub-topics. (2) Dialog Analyzer, which extracts short-term and long-term features from chat history (3) Utterance Strategies Arbitrator, which determines the dialog acts corresponding to our design dimensions, which are outlined in Figure 2. In accordance with the framework, the complete algorithm can be found in Algorithm $1$This algorithm is slightly different from its actual implementation, where $act_{j+1}$ Direct Chatting takes place immediately if a keyword @mubot is matched in the participant’s utterance, see in Appendix A.1. The details of Direct Chatting can be found in Section 3.2.3. Overall, the Sub-topics Generation is executed once, and the Dialog Analyzer and Utterance Strategies Arbitrator are executed sequentially for every $N_{exe}$ utterances, which ensures the latency-efficiency in front of the higher message traffic and complex interactions from multiple users. The following sub-sections provide an in-depth explanation of each module.
@@ 3.2.1 Sub-topic Generator
This module initializes a set of relevant sub-topics, $T$, pretraining to the primary subject in user-input information $I$: $T\sim\bar{p}_{\theta}(T|I)$. User-input information $I$ includes a topic and may optionally include an agenda, hints, and attendee roles. Hints represent any supplementary details users wish to emphasize, while attendee roles refer to the positions designated to participants. The integration of relevant hints and attendee roles has demonstrated success in generating fewer arbitrary and random sub-topics. Consequently, this module enables the MUCA to seamlessly navigate and engage in conversations based on these derived sub-topics. The prompts are documented in Appendix A.2.1.
@@ 3.2.2 Dialog Analyzer
The major task of the Dialog Analyzer is to extract and update useful long-term and short-term signals to help the Utterance Strategies Arbitrator module determine the appropriate dialog acts for the response. It consists of the following sub-modules:
@@
Accumulative Summary Update: This sub-module updates the summary for each participant across various sub-topics based on the long-term context window [^2] Modern LLMs may have extensive context window sizes, with some capable of processing over 32k tokens, which allows chatbots to consider very long historical data, regardless of efficiency and cost considerations. In our experiment, we use a relatively smaller context window $N_{sw}$ to demonstrate that the summarization feature, by design, can be achieved without feeding all chat history, and is also feasible for LLMs with relatively smaller context window sizes. More details see Appendix A.2.2.. It takes into account sub-topics $T_{d}$ (being discussed), previous summary $s_{j}$ and the short-term context window with a size of $N_{sw}$: $s_{j+1}\sim\bar{p}_{\theta}(s_{j+1}|T_{d},s_{j},U_{N_{sw},i})$. The updated summary is subsequently fed back into the current sub-modules (sub-topic status update and accumulative summary update) for further refinements in the next round. Additionally, it is utilized by downstream modules to generate suitable responses. More descriptions are shown in Appendix A.2.2.
Participant Feature Extractor: Unlike the above LLM-based sub-modules, this sub-module extracts statistical features, which include chime-in frequency $freq$, chime-in utterance total length $len$ for each participant from both long-term context window with a size of $N_{lw}$ and short-term context window with a size of $N_{sw}$. Additionally, the proposed participant feature extractor measures the frequency at which each participant chimes in on different sub-topics, effectively documenting their degree of interest in those specific areas. This serves as a reference for customizing encouragement to increase lurkers’ participation – details are discussed in the next sub-section. These statistical data are utilized in a downstream sub-module (participation encouragement) to identify participants who play a passive role in group interactions, and then provide customized encouragement for them to speak up. This sub-module also records the number of participants who discussed the sub-topic from the beginning $N_{ed}$ and the number of participants who are still discussing under the short-term context window $N_{ing}$. These data serve as signals for sub-topic transition in the next sub-section.
@@ 3.2.3 Utterance Strategies Arbitrator
Direct Chatting: Participants can directly interact with MUCA, which serves as a support assistant for individual users, addressing their specific requests and providing assistance as needed. This dialog act always has the highest priority and MUCA responds immediately regardless of the execution period once a user pings the MUCA. Thus, it has no warm-up and cool-down turn, and its trigger condition is met when the keyword of @mubot (case-insensitive) is matched in the last utterance: $cond\leftarrow f(u_{i})$. Many upstream features are extracted by the Dialog Analyzer and used as references for generating appropriate responses: $sig_{j+1}=\{t_{j+1},T_{d},I\}$. It is also worth mentioning that additional well-crafted prompting is required to avoid potential hallucination [^4] When a chatbot is designed based on LLMs, the hallucination is inherited, generally causing confusion and misunderstanding for users. Without careful treatment, the chatbot might provide irrelevant or incorrect information., which is very common especially in this dialog act. Examples can be shown in Section 4.2.2. MUCA demonstrates adaptive responses to various types of requests, shown in Appendix A.2.3.
Initiative Summarization: This dialog act enables the MUCA to generate a concise take-home summary from fragmented messages across various participants and sub-topics, providing a single insightful sentence for a clearer grasp of the discussion. Its trigger condition is heuristically designed for the scenarios when enough participants $N_{active}$ actively joined discussions since the last triggering: $cond\leftarrow f(N_{active},N_{p})$. Its warm-up turn and cool-down turn depend on the number of participants $N_{p}$: $w=11*N_{p}$ and $c=12*N_{p}$, respectively. Accumulative summary update sub-module periodically updates the summary using $sig_{j+1}=\{T_{d}\}$ and concisely presents the key take-home message, which will be displayed when it becomes the highest ranked eligible dialog act.
Participation Encouragement: This dialog act aims to engage less vocal participants and promote balanced contributions in a conversation. It utilizes statistical features derived from participant feature extractor (such as chime-in frequency $freq$, utterance total length $len$) to identify participants who have spoken very little in both long-term and short-term context windows. The process of identifying a participant as a lurker is designed to be conservative. A participant is only considered as a lurker if their $freq$ and $len$ are significantly lower than the average in the long-term context window, and they have spoken very few in the short-term context window, shown in Equation 1 and 2 below. Instead of using KL divergence, which evaluates overall distribution difference, we computed a ratio related to the variance to focus on individual participant data. For conditions based on long-term features:
@@
where $freq^{(x)}$, $mean_{freq}$, $var_{freq}$ in Equation 1 denotes the chime-in frequency for participant $x$, frequency mean and sample variance value for all participants in the long-term window, respectively. $len$ denotes utterance total length. $thre_{freq}=0.4$ and $thre_{len}=0.4$ are chosen heuristically based on experiments. Similarly, $freq^{(x)}$, $len^{(x)}$ in Equation 2 denotes features for participant $x$ in the short-term window and we choose $thre_{freq}=1$ and $thre_{len}=5$.
@@
Also, its warm-up turn is set as $w=3*N_{p}$ and the cool-down turn for each participant increases linearly when triggered each time: $c^{(x)}_{j+1}=2+c^{(x)}_{j}$. This approach ensures that users who prefer to remain silent are not constantly prompted to participate in the chat. MUCA generates personalized messages by considering the lurker’s chat history and chime-in status extracted from participant features: $sig_{j+1}=\{freq_{j+1},len_{j+1}\}$. See more details and examples in the Appendix A.2.3.
@@
Sub-topic Transition: This dialog act enables the MUCA to introduce a new and highly relevant topic when conversations reach a point where users have well-discussed the current topic or when most users are no longer interested in it. There are two trigger conditions for transitioning to a new sub-topic, $cond\leftarrow f(N_{ed},N_{ing},N_{p})$: (1) less than half of the participants have discussed the current subtopic: $N_{ed}<N_{p}/2$, and (2) more than half of the participants were discussing the current subtopic, but only a few are discussing it now within a short-term context window of size $N_{sw}$: $N_{ed}>=N_{p}/2$ and $N_{ing}<N_{ed}/2$. Its warm-up turn and cool-down turn are set as $w=5*N_{p}$ and $c=7*N_{p}$, respectively. Note that its rank is lower than Participation Encouragement since the MUCA aims to encourage inactive participants to contribute before switching to a new sub-topic using $sig_{j+1}=\{N_{ed},N_{ing}\}$. Introducing a new sub-topic may disrupt the conversation flow and potentially divert the discussion from its current trajectory. Examples are shown in the Appendix A.2.3.
Figure 4: Illustration of user simulator (MUS).
@@
Conflict Resolution: This dialog act helps participants reach a consensus in a timely manner, thereby providing an efficient discussion procedure. Different from previous study which set time limitations for each task Kim et al. (2020a), MUCA provides suggestions to help parties with diverse opinions reach a consensus, while also suggesting the next topic for discussion. (See the example in Section 4.2.2). We utilize a similar template as sub-topic transition, but with an added requirement of a summary to facilitate suggestion generation. Its trigger condition is met when the number of well-discussed sub-topics does not increase for a given period of time $=9*N_{p}$: $cond\leftarrow f(N_{well},9*N_{p})$. Its warm-up turn is set as $w=9*N_{p}$ and shares the same cool-down and same $sig_{j+1}$ with sub-topic transition in our design.
@@
In-context Chime-in: The MUCA features an automatic chime-in dialog act designed to enhance conversation depth by providing insights, advancing stuck chatting scenarios, and addressing participants’ concerns. Its trigger condition is $cond\leftarrow p_{chime-in}>thre$, where $thre=0.45$ and two factors contributing to the in-context chime-in probability $p_{chime-in}$: (1) silence factor probability $p_{silence}$: it increases with the number of consecutive silent turns $n_{silent}$ (See Equation 3). Thus, the chatbot has a higher probability to chime-in when it has been silent for multiple turns. (2) semantic factor probability $p_{semantic}$ (See Equation 4): it is associated with situations where the conversation is stuck due to repetitive utterances or unresolved issues that the chatbot must address. Questions and concerns exchanged between participants do not constitute unsolved issues for the chatbot. Appropriate CoT-style prompting is used to identify if the conversation is stagnant. $p_{chime-in}$ is calculated by averaging the $p_{silence}$ and $p_{semantic}$ (See Equation 5). It has no additional $w$ and $c$, and it uses the same $sig_{j+1}$ as direct chatting since it also needs to provide information that requires the long-term context beyond the short-term context window. Examples are shown in the Appendix A.2.3.
@@
Keep Silence: This dialog act allows the MUCA to refrain from excessive responses, preventing irritation in a multi-user setting. Keep Silence is automatically activated when other dialog acts’ triggering conditions are not met, maintaining the conversation’s flow without distracting participants.
@@ 3.3 User Simulation
Figure 5: Log-normal distribution, which is clipped at a minimum length of utterance $l_{min}=3$ and a maximum length of utterance $l_{max}=20$.
@@ 3.3.1 MUS Framework Design
The proposed MUS, which consists of two primary modules, is depicted in Figure 4. All implementations are based on well-crafted prompting of LLMs. User Behavior Resembling module consumes previous group conversation snippets $C_{s}$ to generate user behavior signals in terms of speaking role $S_{r}^{(x)}$, utterance traits $U_{t}^{(x)}$ and utterance length $l_{utt}^{(x)}$. This module only executes once right before the simulation starts, as shown in Algorithm 2. User Utterance Generation module utilizes these signals to produce natural language utterances $utt_{i+1}$, which mimics real user behavior. It executes $N_{exe}$ times to generate utterances and then MUCA generates a corresponding response based on the context. More details for each module are shown below.
User Behavior Resembling: Initially, three chat snippets ($C_{s}$) are chosen from actual user conversation logs, containing 10-30 turns each. To define the utterance content and idiolect, 11 speaking roles ($S_{r}$) and 10 utterance traits ($U_{t}$) are predetermined. For example, questioner, as a speaking role, refers to someone who raises a question based on the given context, while laconic, as an utterance trait, denotes the expression of thoughts using the fewest words possible. We select one subset of roles and one subset of traits for each virtual attendee: $S_{r}^{(x)},U_{t}^{(x)}\sim\bar{p}_{\theta}^{CoT}(S_{r}^{(x)},U_{t}^{(x)}|C_{s})$, where $S_{r}^{(x)}\subset S_{r}$ and $U_{t}^{(x)}\subset U_{t}$. Each subset $S_{r}^{(x)}$ and $U_{t}^{(x)}$ contains $M$ ($=6$) speaking roles and utterance traits respectively, which are learned from chat snippets via LLM prompting, which is shown in the Appendix A.2.4.
We also simulate a log-normal distribution for the utterance length (i.e., word count) of each virtual attendee, using the minimum length $l_{min}$, maximum length $l_{max}$ and average length $l_{avg}$ derived from a 150-turn group chat. This is demonstrated in Equation 6, 7 and 8 below. The distribution is skewed towards smaller lengths and is applied in the User Utterance Generation module. Figure 5 shows an example of log-normal distribution for utterance length $l_{utt}$. The User Behavior Resembling module is tailored to each virtual attendee and initialized offline. This module governs the content and idiolect of utterances to maintain consistent behavior among virtual attendees by utilizing the historical data from real users as a reference.
@@
where $\mu$ is the mean and $\sigma$ is the standard deviation of the normally distributed logarithm of the variable. The estimation of $\mu$ and $\sigma$ remains heuristic and uses minimum length $l_{min}$, maximum length $l_{max}$ and average length $l_{avg}$, where $w=0.3$ and $\alpha=0.67$.
User Utterance Generation: Given the conversation context (=16 turns), chat history summary from LLM agent (see in Section 3.2), CoT-style prompting is applied here to firstly generate a virtual attendee for the next turn: $v_{i+1}\sim\bar{p}_{\theta}^{CoT}(v_{i+1}|I,U_{N_{sw},i})$ and then the corresponding proper speaking role: $sr_{i+1}\sim\bar{p}_{\theta}^{CoT}(sr_{i+1}|v_{i+1},U_{N_{sw},i})$. In our instruction, we implement a constraint ensuring that the subsequent virtual participant is neither the same individual from the last turn nor the LLM agent. It is worth mentioning that GPT-4 may not consistently follow the given instructions, particularly when it generates the LLM agent in response to being directly addressed in the previous turn. In this scenario, the virtual attendee and their speaking role are randomly selected. Following this, the utterance length for the chosen attendee is randomly determined from the log-normal distribution: $l_{utt}\sim\text{LogNorm}(x|\mu,\sigma^{2})$. The CoT approach is then applied once more to initially generate the utterance trait, which is affected by the utterance length and other utterance features (previous summary): $ut_{i+1}\sim\bar{p}_{\theta}^{CoT}(ut_{i+1}|l_{utt},v_{i+1},s_{j+1},I,C_{s},sr_{i+1},U_{N_{sw},i})$. Subsequently, the utterance is produced, taking into account all generated signals mentioned above, see Algorithm 2. Limitations and ad-hoc solutions are also discussed in Appendix 5.1. See prompting details and examples in Appendix A.2.4.

### 16 Feb 2024 v3

@@ Abstract
Recent advancements in large language models (LLMs) have provided a new avenue for chatbot development, while most existing research has primarily centered on single-user chatbots that focus on deciding "What" to answer after user inputs. In this paper, we identified that multi-user chatbots have more complex 3W design dimensions – "What" to say, "When" to respond, and "Who" to answer. Additionally, we proposed Multi-User Chat Assistant (MUCA), which is an LLM-based framework for chatbots specifically designed for group discussions. MUCA consists of three main modules: Sub-topic Generator, Dialog Analyzer, and Utterance Strategies Arbitrator. These modules jointly determine suitable response contents, timings, and the appropriate recipients. To make the optimizing process for MUCA easier, we further propose an LLM-based Multi-User Simulator (MUS) that can mimic real user behavior. This enables faster simulation of a conversation between the chatbot and simulated users, making the early development of the chatbot framework much more efficient. MUCA demonstrates effectiveness, including appropriate chime-in timing, relevant content, and improving user engagement, in group conversations with a small to medium number of participants, as evidenced by case studies and experimental results from user studies.
@@ 2 Related Work
LLMs have demonstrated remarkable performance across various tasks. They are typically pre-trained on large amounts of data and fine-tuned with human-curated data through supervised learning and reinforcement learning approaches. Some of the recent work including GPT models OpenAI (2023); Ouyang et al. (2022); Brown et al. (2020); Radford et al. (2019b), Gemini Team et al. (2023), LLaMa2 Touvron et al. (2023), PaLM2 Anil et al. (2023), Bloom Scao et al. (2022) and LaMDa Thoppilan et al. (2022). Some of the recent versions of these models, such as GPT-4, have shown promising results as conversational agents. Despite the overall good performance of these models across multiple tasks on text generation, there are still continuing efforts to improve them for specific scenarios Chen et al. (2021); Wu et al. (2023b).
@@ 3.2 Framework Architecture
Figure 3 illustrates the framework architecture and information flow, which consists of three major modules: (1) Sub-topics Generator, which generates the initial sub-topics. (2) Dialog Analyzer, which extracts short-term and long-term features from chat history (3) Utterance Strategies Arbitrator, which determines the dialog acts corresponding to our design dimensions, which are outlined in Figure 2. In accordance with the framework, the complete algorithm can be found in Algorithm $1$This algorithm is slightly different from its actual implementation, where $act_{j+1}$ Direct Chatting takes place immediately if a keyword @mubot is matched in the participant’s utterance, see in Appendix A.1. The details of Direct Chatting can be found in Section 3.2.3. Overall, the Sub-topics Generation is executed once, and the Dialog Analyzer and Utterance Strategies Arbitrator are executed sequentially for every $N_{exe}$ utterance, which ensures the latency-efficiency in front of the higher message traffic and complex interactions from multiple users. The following sub-sections provide an in-depth explanation of each module.
@@ 3.2.1 Sub-topic Generator
This module initializes a set of relevant sub-topics, $T$, from the primary subject in user-input information $I$: $T\sim\bar{p}_{\theta}(T|I)$. User-input information $I$ includes a topic and may optionally include an agenda, hints, and attendee roles. Hints represent any supplementary details users wish to emphasize, while attendee roles refer to the positions designated to participants. The integration of relevant hints and attendee roles has demonstrated success in generating fewer arbitrary and random sub-topics. Consequently, this module enables the MUCA to seamlessly navigate and engage in conversations based on these derived sub-topics. The prompts are documented in Appendix A.2.1.
@@ 3.2.2 Dialog Analyzer
This module extracts useful long-term and short-term signals from the conversation that are deemed useful to MUCA to understand the ongoing conversation in order to decide the best action to take, and therefore it is a constantly executing module that starts running once the chat begins. It consists of the following sub-modules:
@@
Accumulative Summary Update: This sub-module updates the summary for each participant under different sub-topics based on the all chat history [^2] Modern LLMs may have extensive context window sizes, with some capable of processing over 32k tokens, which allows chatbots to consider very long historical data, regardless of efficiency and cost considerations. In our experiment, we use a relatively smaller context window $N_{sw}$ to demonstrate that the summarization feature, by design, can be achieved without feeding all chat history, and is also feasible for LLMs with relatively smaller context window sizes. More details see Appendix A.2.2. by considering sub-topics $T_{d}$, previous summary $s_{j}$ and $U_{N_{sw},i}$: $s_{j+1}\sim\bar{p}_{\theta}(s_{j+1}|T_{d},s_{j},U_{N_{sw},i})$. This essentially builds a memory into the MUCA system. The updated summary is subsequently fed back into the current sub-modules (sub-topic status update and accumulative summary update) for further refinements in the next round. More descriptions are shown in Appendix A.2.2.
Participant Feature Extractor: Unlike the above LLM-based sub-modules, this sub-module extracts statistical features, which include chime-in frequency $freq$, chime-in utterance total length $len$ for each participant from both long-term context window with a size of $N_{lw}$ and short-term context window with a size of $N_{sw}$. Additionally, this sub-module also measures the frequency at which each participant chimes in on different sub-topics, effectively documenting their degree of interest in those specific areas. This serves as a reference for customizing encouragement to increase lurkers’ participation. The number of participants who discussed the sub-topic from the beginning $N_{ed}$ and the number of participants who are still discussing under the short-term context window $N_{ing}$ serve as signals for sub-topic transition in Utterance Strategies Arbitrator.
@@ 3.2.3 Utterance Strategies Arbitrator
Direct Chatting: Participants can directly interact with MUCA, which serves as a support assistant for individual users, addressing their specific requests and providing assistance as needed. This dialog act always has the highest priority and MUCA responds immediately regardless of the execution period once a user pings the MUCA. Thus, it has no warm-up and cool-down turn, and its trigger condition is met when the keyword of @mubot (case-insensitive) is matched in the last utterance: $cond\leftarrow f(u_{i})$. Many upstream features are extracted by the Dialog Analyzer and used as references for generating appropriate responses: $sig_{j+1}=\{t_{j+1},T_{d},I\}$. It is also worth mentioning that additional well-crafted prompting is required to avoid potential hallucination [^4] When a chatbot is designed based on LLMs, hallucination issues may be inherited, generally causing confusion and misunderstanding for users. Without careful treatment, the chatbot might provide irrelevant or incorrect information., which is very common especially in this dialog act. Examples can be found in Section 4.2.2. MUCA demonstrates adaptive responses to various types of requests, shown in Appendix A.2.3.
Initiative Summarization: This dialog act enables the MUCA to generate a concise take-home summary from fragmented messages across various participants and sub-topics, providing a single insightful sentence for a clearer grasp of the discussion. Its trigger condition is heuristically designed for the scenarios when enough participants $N_{active}$ actively joined discussions since the last triggering: $cond\leftarrow f(N_{active},N_{p})$. Its warm-up turn and cool-down turn depend on the number of participants $N_{p}$: $w=11*N_{p}$ and $c=12*N_{p}$, respectively. Accumulative summary update sub-module periodically updates the summary using $sig_{j+1}=\{T_{d}\}$ and concisely presents the key take-home message, which will be displayed when Initiative Summarization becomes the highest ranked eligible dialog act. Initiative Summarization uses the same prompting template of Accumulative summary update but it has own warm-up turn and cool-down turn.
Participation Encouragement: This dialog act utilizes statistical features derived from participant feature extractor (such as chime-in frequency $freq$, utterance total length $len$) to identify participants who have spoken very little in both $U_{N_{sw},i}$ and $U_{N_{lw},i}$. The process of identifying a participant as a lurker is designed to be conservative. A participant is only considered as a lurker if their $freq$ and $len$ are significantly lower than the average in the long-term context window, and they have spoken very little in the $U_{N_{sw},i}$, as determined by Equation 1 and 2 below. Instead of using measures like KL divergence which evaluates overall distribution difference, we compute a ratio related to the variance to focus on individual participant data. For conditions based on long-term features:
@@
where $freq^{(x)}$, $mean_{freq}$, $var_{freq}$ in Equation 1 denotes the chime-in frequency for participant $x$, frequency sample mean and frequency sample variance value for all participants in the long-term window, respectively. $len$ denotes utterance total length in word counts. $thre_{freq}=0.4$ and $thre_{len}=0.4$ are chosen heuristically based on experiments. Similarly, $freq^{(x)}$ and $len^{(x)}$ in Equation 2 denote features for participant $x$ in the short-term window and we choose $thre_{freq}=1$ and $thre_{len}=5$.
@@
The warm-up turn is set to $w=3*N_{p}$ and the cool-down turn for each participant increases linearly after each trigger: $c^{(x)}_{j+1}=2+c^{(x)}_{j}$. This approach ensures that users who prefer to remain silent are not constantly pinged to participate in the chat. MUCA generates personalized messages by considering the customized inactive status of the lurker, which is generated from the lurker’s chat history and participation features: $sig_{j+1}=\{freq_{j+1},len_{j+1}\}$. See more details and examples in the Appendix A.2.3.
@@
Sub-topic Transition: This dialog act enables the MUCA to introduce a new and highly relevant topic when conversations reach a point where users have well-discussed the current topic or when most users are no longer interested in it. So it has two trigger conditions, both of which take the form of $cond\leftarrow f(N_{ed},N_{ing},N_{p})$. For condition 1, the triggering criterion is when less than half of the participants have discussed the current subtopic: $N_{ed}<N_{p}/2$. For condition 2, the triggering criterion is when more than half of the participants were discussing the current subtopic, but only a few are discussing it now within $U_{N_{sw},i}$: $N_{ed}>=N_{p}/2$ and $N_{ing}<N_{ed}/2$. The warm-up turn and cool-down turn are set as $w=5*N_{p}$ and $c=7*N_{p}$, respectively. Note that its priority is lower than Participation Encouragement since MUCA encourages lurkers to contribute before considering whether to start a new sub-topic using $sig_{j+1}=\{N_{ed},N_{ing}\}$. Introducing a new sub-topic may disrupt the conversation flow and potentially divert the discussion from its current status. Examples are shown in the Appendix A.2.3.
Figure 4: Illustration of multi-user simulator (MUS).
@@
Conflict Resolution: This dialog act helps participants reach a consensus in a timely manner, thereby improving the efficiency of the conversation. Different from previous studies which set time limitations for each task Kim et al. (2020a), MUCA provides suggestions to help parties with diverse opinions reach a consensus, and at the same time suggests a next topic for discussion. An example is given in Sec. 4.2.2. We utilize a prompt template somewhat similar to that of sub-topic transition, but with an added requirement of including a summary in its response to facilitate suggestion generation. Its trigger condition is met when the number of well-discussed sub-topics does not increase for a given period of time $=9*N_{p}$: $cond\leftarrow f(N_{well},9*N_{p})$. Its warm-up turn is set as $w=9*N_{p}$ and shares the same cool-down and same $sig_{j+1}$ with sub-topic transition in our design.
@@
In-context Chime-in: MUCA features an automatic chime-in dialog act designed to enhance conversation depth by providing insights, advancing stuck chatting scenarios, and addressing participants’ concerns. Its trigger condition is $cond\leftarrow p_{chime-in}>thre$, where $thre=0.45$ and two factors contributing to the in-context chime-in probability $p_{chime-in}$: (1) silence factor probability $p_{silence}$: it increases with the number of consecutive silent turns $n_{silent}$ (See Equation 3). Thus, the chatbot has a higher probability to chime-in when it has been silent for multiple turns. (2) semantic factor probability $p_{semantic}$ (See Equation 4): it is associated with situations where the conversation is stuck due to repetitive utterances or unresolved issues that the chatbot must address. Questions and concerns exchanged between participants do not constitute unsolved issues for the chatbot. Appropriate CoT-style prompting is used to identify if the conversation is stagnant. $p_{chime-in}$ is calculated by averaging the $p_{silence}$ and $p_{semantic}$ (See Equation 5). It has no additional $w$ and $c$, and it uses the same $sig_{j+1}$ as direct chatting since it also needs to provide information that requires the long-term context beyond the $U_{N_{sw},i}$. Examples are shown in the Appendix A.2.3.
@@
Keep Silent: This dialog act allows the MUCA to refrain from excessive responses, preventing irritation in a multi-user setting. Keep Silent is automatically activated when other dialog acts’ triggering conditions are not met, maintaining the conversation’s flow without distracting participants.
@@ 3.3 User Simulation
Figure 5: (Discretized and truncated) log-normal distribution, which is clipped at a minimum length of utterance $l_{min}=3$ and a maximum length of utterance $l_{max}=20$.
@@ 3.3.1 MUS Framework Design
The proposed MUS, which consists of two primary modules, is depicted in Figure 4. Implementations are largely based on prompt engineering and LLMs. User Behavior Resembling module consumes previous group conversation snippets $C_{s}$ to generate user behavior signals in terms of speaking role $S_{r}^{(x)}$, utterance traits $U_{t}^{(x)}$ and utterance length $l_{utt}^{(x)}$. This module only executes once right before the simulation starts, as shown in Algorithm 2. User Utterance Generation module utilizes these signals to produce natural language utterances $utt_{i+1}$, which mimics real user behavior. It executes $N_{exe}$ times to generate utterances and then MUCA generates a corresponding response based on the context. More details for each module are shown below.
User Behavior Resembling: Initially, three chat snippets ($C_{s}$) are chosen from actual user conversation logs, containing 10-30 turns each. To define the utterance content and idiolect, 11 speaking roles ($S_{r}$) and 10 utterance traits ($U_{t}$) are predetermined. For example, questioner, as a speaking role, refers to someone who raises a question based on the given context, while laconic, as an utterance trait, denotes the expression of thoughts using the fewest words possible. We select one subset of roles and one subset of traits for each virtual attendee: $S_{r}^{(x)},U_{t}^{(x)}\sim\bar{p}_{\theta}^{CoT}(S_{r}^{(x)},U_{t}^{(x)}|C_{s})$, where $S_{r}^{(x)}\subset S_{r}$ and $U_{t}^{(x)}\subset U_{t}$. Each subset $S_{r}^{(x)}$ and $U_{t}^{(x)}$ contains $M$ ($=6$) speaking roles and utterance traits respectively, which are learned from chat snippets via LLM prompting, which is shown in the Appendix A.3.
We sample from a log-normal distribution (see Equation 6, 7 and 8) for the utterance length (i.e., word count) of each virtual attendee. The distribution was constructed using the minimum length $l_{min}$, maximum length $l_{max}$ and average length $l_{avg}$ derived from a 150-turn group chat, and is also discretized and truncated. The distribution is skewed towards smaller lengths and is applied in the User Utterance Generation module. Figure 5 shows an example of log-normal distribution for utterance length $l_{utt}$. The User Behavior Resembling module is tailored to each virtual attendee and initialized offline. This module governs the content and idiolect of utterances to maintain consistent behavior among virtual attendees by utilizing the historical data from real users as a reference.
@@
where $\mu$ is the mean and $\sigma$ is the standard deviation of the normally distributed logarithm of the variable. In this work, we use $w=0.3$ and $\alpha=0.67$.
User Utterance Generation: Given the conversation context gathered from most recent 16 turns, chat history summary from LLM agent (see in Section 3.2), CoT-style prompting is applied here to first determine a virtual attendee for the next turn: $v_{i+1}\sim\bar{p}_{\theta}^{CoT}(v_{i+1}|I,U_{N_{sw},i})$. The module then determines the corresponding proper speaking role: $sr_{i+1}\sim\bar{p}_{\theta}^{CoT}(sr_{i+1}|v_{i+1},U_{N_{sw},i})$. In our instruction, we implement a requirement to guide the module so that the subsequent virtual participant is neither the same individual from the last turn nor the LLM agent. It is worth mentioning that an LLM is not guaranteed to consistently follow the given instructions, particularly when it generates the LLM agent in response to being directly addressed in the previous turn. In this scenario, the virtual attendee and their speaking role are randomly selected. Following this, the utterance length for the chosen attendee is randomly determined from the log-normal distribution: $l_{utt}\sim\text{LogNorm}(x|\mu,\sigma^{2})$. The CoT approach is then applied once more to initially generate the utterance trait, which is affected by the utterance length and other utterance features (previous summary): $ut_{i+1}\sim\bar{p}_{\theta}^{CoT}(ut_{i+1}|l_{utt},v_{i+1},s_{j+1},I,C_{s},sr_{i+1},U_{N_{sw},i})$. Subsequently, the utterance is produced, taking into account all generated signals mentioned above, see Algorithm 2. Limitations and ad-hoc solutions are discussed in Section 5. See prompting details and examples in Appendix A.3.